# SimpleDeepSearcher: Deep Information Seeking via Web-Powered Reasoning Trajectory Synthesis

### v2

@@ 4.3 Response Redundancy
4.4 Stage-wise Analysis
In this part, we analyze how training improves the performance of each sub-task in our approach, including iterative search, planning, and summarization. We evaluate the proportion of cases in which the final answer appears during each sub-process to quantify the efficiency of that stage. To eliminate interference from the summarization stage, all summarization models are kept identical during inference, with detailed settings provided in Appendix D. The results are shown in Table  5. We can observe substantial improvements across all components, with the search component showing the most significant gain. This suggests that training effectively enhances the model’s ability to generate more coherent reasoning and search trajectories, leading to more accurate information retrieval and improved overall model performance.
Models Summarization Models Bamboogle GAIA
@@ 4.4 Stage-wise Analysis
Table 6: Performance comparison across two benchmarks using different summarization models.
@@ 4.5 Effect of Summarization Model
This part investigates the impact of the summarization model on overall performance. We fix the reasoning model and conduct a comparative analysis of overall performance using different summarization models. As shown in Table 6, QwQ-32B demonstrates the strongest summarization capability and is therefore selected as the summarization model for all reasoning models. Furthermore, using fine-tuned models for summarization leads to performance degradation on downstream tasks compared to their pre-trained counterparts. This might be attributed to the reduced long-text summarization ability of the fine-tuned models, due to the distributional shifts on a limited task and domain of the training data. This decline is more pronounced for models with fewer parameters.
@@ 4.6 Effect of Additional Reasoning Data
Figure 2: Average reasoning length across three benchmarks w/ and w/o reasoning data for training.
@@
Table 8: Average search count across three benchmarks of the model trained w/ and w/o reasoning data.
@@ Appendix A Related Work
Enhancing LLMs with Search. Recently, several deep search frameworks are proposed (Ren et al., 2025; Alzubi et al., 2025). They integrate large language models with search engines in a more flexible and dynamic manner. Search-o1 (Li et al., 2025a) simulates deep search in LLMs through prompt engineering, allowing them to retrieve information independently during multi-step reasoning. R1-Searcher (Song et al., 2025) and Search-R1 (Jin et al., 2025) equip large language models with retrieval tools and train them end-to-end using reinforcement learning. This approach effectively enhances the model’s ability to interleave reasoning with retrieval during inference. However, due to the inherent complexity of RL and its high computational demands, conducting large-scale experiments on full-sized LLMs remains challenging. SimpleDeepSearcher synthesizes high-quality training data via broad query sampling and precise filtering, enabling strong deep search performance with minimal training cost.
Appendix B DPO Detailed Settings
Our objective was to identify answer trajectories that were both correct and demonstrated efficient reasoning and search paths. To this end, we construct preference pairs $(R_{w},R_{l})$, where $R_{w}$ denotes the preferred trajectory and $R_{l}$ the rejected one. We repurpose our previously established pipeline for query sampling and data synthesis. During the data synthesis stage, we generate responses using the strongest SFT-trained model, SimpleDeepSearcher-QwQ-32B-SFT, and the target model to be optimized, SimpleDeepSearcher-Qwen-7B-SFT. Responses generated by SimpleDeepSearcher-QwQ-32B-SFT that pass both the formatting and reasoning path control checks are treated as chosen examples, while those generated by SimpleDeepSearcher-Qwen-7B-SFT that fail these checks are treated as rejected examples. Ultimately, we construct a dataset consisting of approximately 875 training pairs.
@@ Appendix B DPO Detailed Settings
Appendix C REINFORCE++ Detailed Settings
@@
The reward function employed in REINFORCE++ consists of two components: an answer reward and a format penalty. The answer reward is calculated as the F1 score between the predicted answer and the reference answer, providing a direct measure of response accuracy. In addition, a discrete format penalty of $-2$ is applied if any of the following undesirable behaviors are detected:
@@
Appendix D Stage-wise Analysis Settings
We conduct a comparative analysis of Qwe2.5-7B-Instruct and QwQ-32B before and after training across the 2Wiki, MuSiQue, and Bamboogle benchmarks. During inference, we fix the summarization model to QwQ-32B across all comparisons to eliminate potential interference from the summarization component (the impact of the summarization model will be further discussed in Section 4.5).
Appendix E Instruction Templates
Figure 4 shows the instruction for annotating the domain and keywords of questions. Figure 5 shows the instruction for LLM as a judge. Figure 6 shows the instruction for the reasoning model. Figure 7 shows the instruction for the summarization model.
Figure 4: Instruction for annotation.
Figure 5: Instruction for LLM as a judge.
Figure 6: Instruction for reasoning model.
Figure 7: Instruction for summarization model.

### v3

@@ 4.3 Effect of SFT on Redundancy
4.4 Effect of SFT on Stage-wise Performance
In this part, we analyze how training improves the performance of each sub-task in our approach, including iterative search, planning, and summarization. We evaluate the proportion of cases in which the final answer appears during each sub-process to quantify the efficiency of that stage. To eliminate interference from the summarization stage, all summarization models are kept identical during inference, with detailed settings provided in Appendix E. The results are shown in Table  6. We can observe substantial improvements across all components, with the search component showing the most significant gain. This suggests that training effectively enhances the model’s ability to generate more coherent reasoning and search trajectories, leading to more accurate information retrieval and improved overall model performance.
Models Summarization Model Bamboogle GAIA
@@ 4.4 Effect of SFT on Stage-wise Performance
Table 7: Performance comparison across all benchmarks using different summarization models.
@@ 4.5 Effect of Summarization Model
This part investigates the impact of the summarization model on overall performance. We fix the reasoning model and conduct a comparative analysis of overall performance using different summarization models. As shown in Table  7, QwQ-32B demonstrates the strongest summarization capability and is therefore selected as the summarization model for all reasoning models. Furthermore, using fine-tuned models for summarization leads to performance degradation on downstream tasks compared to their pre-trained counterparts. This might be attributed to the reduced long-text summarization ability of the fine-tuned models, due to the distributional shifts on a limited task and domain of the training data. This decline is more pronounced for models with fewer parameters.
@@ 4.6 Effect of Additional Reasoning Data
Figure 2: Average reasoning length across different benchmarks w/ and w/o reasoning data for training.
@@
Table 9: Average search count across different benchmarks of the model trained w/ and w/o reasoning data.
@@ 5 Conclusion
Ethics Statement
We strictly adhere to ethical standards. We follow the relevant licenses and guidelines for dataset usage, ensuring that no personal or offensive information is included. AI assistance was only utilized during the paper refinement process. Our trained models do not display any potential biases or discriminatory behavior, and we rigorously comply with research ethics throughout the entire development and evaluation process.
Acknowledgments
This work was partially supported by National Natural Science Foundation of China under Grant No. 92470205 and 62222215, Beijing Natural Science Foundation under Grant No. L233008 and Beijing Municipal Science and Technology Project under Grant No. Z231100010323009.
@@ Appendix A Related Work
Enhancing LLMs with Search. Recently, several deep search frameworks are proposed (Alzubi et al., 2025). They integrate large language models with search engines in a more flexible and dynamic manner. Search-o1 (Li et al., 2025a) simulates deep search in LLMs through prompt engineering, allowing them to retrieve information independently during multi-step reasoning. R1-Searcher (Song et al., 2025) and Search-R1 (Jin et al., 2025) equip large language models with retrieval tools and train them end-to-end using reinforcement learning. This approach effectively enhances the model’s ability to interleave reasoning with retrieval during inference. However, due to the inherent complexity of RL and its high computational demands, conducting large-scale experiments on full-sized LLMs remains challenging. SimpleDeepSearcher synthesizes high-quality training data via broad query sampling and precise filtering, enabling strong deep search performance with minimal training cost.
Appendix B Details of Diversity-Aware Query Sampling
In analyzing open-source data, we identified three critical limitations:
(1) Domain-specific overrepresentation creating skewed knowledge distributions. As shown in the Figure 3, we present the domain distribution of the pre-filtered data. It can be observed that certain domains (such as film and geography) account for a considerable proportion. This imbalance risks inducing uneven knowledge distributions in the training data.
(2) Repetitive syntactic patterns reducing linguistic variability. Due to the construction methods of open-source datasets, we observe substantial redundancy in syntactic structures. A typical example is the prevalence of “A and B” style comparative queries (e.g., “Do both directors of films Paper Bullets and Karakolda Ayna Var share the same nationality?” vs. “Do both directors of films Jatt Juliet and Sciopèn share the same nationality?”). Similarly, numerous queries repetitively compare identical attributes such as age.
(3) Semantic simplicity thresholds below real-world query complexity. Many queries in open-source datasets are overly simplistic, such as “What nationality is John Harbaugh’s father?”. Such questions impose only minimal demands on deep search or reasoning, as they can often be answered through a single lookup. Consequently, their utility in fostering more advanced model capabilities is limited.
We define “core semantic constituents” as follows:
$\bullet$ Key entities (e.g., films, people, locations)
$\bullet$ Critical attributes (e.g., age, duration, population)
$\bullet$ Core relationships (e.g., comparison, causality)
$\bullet$ Measurement dimensions (e.g., time, quantity)
For example, in the query “Which film whose director is younger, Charge It To Me or Danger: Diabolik?”, the extracted keywords based on the above schema are “film” and “age”.
Figure 3: Domain distribution of the data before filtering.
Appendix C DPO Detailed Settings
Our objective was to identify answer trajectories that were both correct and demonstrated efficient reasoning and search paths. To this end, we construct preference pairs $(R_{w},R_{l})$, where $R_{w}$ denotes the preferred trajectory and $R_{l}$ the rejected one. We repurposed our previously established pipeline for query sampling and data synthesis. During the data synthesis stage, we generate responses using the strongest SFT-trained model, SDS-QwQ-32B-SFT, and the target model to be optimized, SDS-Qwen-7B-SFT. Responses generated by SDS-QwQ-32B-SFT that pass both the formatting and reasoning path control checks are treated as positive examples, while those generated by SDS-Qwen-7B-SFT that fail these checks are treated as negative examples. Ultimately, we construct a dataset consisting of approximately 875 training pairs.
@@ Appendix C DPO Detailed Settings
Appendix D REINFORCE++ Detailed Settings
@@
The reward function employed in REINFORCE++ consists of two components: an answer reward and a format penalty. The answer reward is calculated as the F1 score between the predicted answer and the reference answer, providing a direct measure of response accuracy. In addition, a discrete format penalty of 2-2 is applied if any of the following undesirable behaviors are detected:
@@
Appendix E Model Performance Enhancement Analysis Settings
We conduct a comparative analysis of Qwe2.5-7B-Instruct and QwQ-32B before and after training across the 2Wiki, Bamboogle, and MuSiQue benchmarks. During inference, we fix the summarization model to QwQ-32B across all comparisons to eliminate potential interference from the summarization component (the impact of the summarization model will be further discussed in Section 4.5).
Appendix F Instruction Templates